OpenStudy (anonymous):

solve the equation using the square root method: 5(x+1)²-2=13

OpenStudy (anonymous):

also how do you get the answer

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

i got to 5x+3=+/- √13

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

5x+5^2-2=13

OpenStudy (anonymous):

exponents first

OpenStudy (anonymous):

5x+25-2=13

OpenStudy (anonymous):

i have to use the square root method

OpenStudy (anonymous):

either method will get you same answer to x

OpenStudy (anonymous):

yes but i need to know how to use the square root method

OpenStudy (anonymous):

oh idk then

OpenStudy (anonymous):

\[5(x+1)^{2} - 2 = 13\] \[=> 5(x+1)^{2}=15\] Now dividing by 5 we get \[(x+1)^{2} =3 \] \[=> x + 1 = \pm \sqrt{3}\] \[\therefore x = \pm \sqrt{3} - 1 \]

OpenStudy (anonymous):

Einstein thanked all the people who said him no when he asked him questions , for it helped his discover what we read and study today. So probably you should curse me fir it , and pledge to work on your own in near future .

OpenStudy (anonymous):

please be quiet